Physics II SP2019 KH01 Problem 14.5-Enhanced- with Feedback Review Part A A 1.60 m -diameter vat of liquid is 2.50 m deep. The pressure at the bottom of the vat is 1.30 atm You may want to review (Pages 360-364) What is the mass of the liquid in the vat? Express your answer with the appropriate units. alue Provide Feedback

A 1.60 m -diameter vat of liquid is 2.50 m deep. The pressure at the bottom of the vat is 1.30 atm. You may want to review (Pages 360-364) 


Part A

What is the mass of the liquid in the vat? Express your answer with the appropriate units. 
Answer #1

Then i took the height of the cyclinder to just bethe height of the liquid as given as 2.50 m

Gauge pressure=1.3-1=0.3 atm=0.3*10^5 Pa

p=pgh

0.3*10^5 = p * 9.81 * 2.5

p=1223.5 kgm^(-3)

p=m/v

1223.5=m/(pi * 0.8^2 * 2.5)

m=6146.86 kg Answer

I hope help you !!